Grade 12Physical Chemistry

A sample of non stoichiometric ferrous oxide has 12% of Fe+3 and remain Fe+2 .what is formula of the compound.

Anjali
9 Years agoGrade 12

1 Answer

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
Let the formule be FexO1.0
And Fe2+ proportion be y then Fe3= = > (x – y)
charge on O = charges on Fe
Therfore,
2y + 3(x-y) = 2*1
solving we get,
y = 3x – 2  = 0.82
x = 0.94
Hence the formulae is: =>
Fe0.94O1.0